OpenStudy (anonymous):

A quadratic function is defined by f(x) = 3x^2 + 4x - 2. A linear function is defined by g(x) = mx-5. What values of the slope of the line would make it a tangent to the parabola?

OpenStudy (anonymous):

After I equated the function, I got 3x^2 + 4x - mx + 3 = 0. b^2 -4ac = 0 Is b = 4-m?

OpenStudy (anonymous):

heelp D:

OpenStudy (anonymous):

The answer is supposed to be m = 10, m = -2

sam (.sam.):

3x^2 + 4x - 2= mx-5 3x^2 + (4-m)x + 3=0 a=3 b=4-m c=3 b^2 -4ac = 0

OpenStudy (anonymous):

why is it (4-m)x?

OpenStudy (anonymous):

that's the thing that's confusing me..

sam (.sam.):

3x^2 + 4x - 2= mx-5 3x^2 + 4x - 2 -mx +5=0 3x^2 + 4x -mx + 3= 0 3x^2 + x(4-m) + 3=0

sam (.sam.):

common factor 4x -mx x(4-m)

OpenStudy (anonymous):

ohhh you factored it. Thank you! :')

OpenStudy (anonymous):

so I don't put b^2 - 4ac = x(4-m)^2 - 4(3)(3)?

sam (.sam.):

without the x

OpenStudy (anonymous):

why not?

sam (.sam.):

(4-m)^2 - 4(3)(3)=0

OpenStudy (amistre64):

you want the derivative of the quadratic to have the same slope as the linear equation and pass thru the point 0,5

OpenStudy (anonymous):

what is the linear equation's slope?

OpenStudy (amistre64):

y = f'(0)x - 5 is the equation of the line that you want to equate f(x) = 3x^2 + 4x - 2 f'(x) = 6x +4 f'(0) = 0+4 = 4 y = 4x - 5 ; so the points on the parabola with a slope of 4

OpenStudy (amistre64):

ideally there should be 2 values for the slope of this line

OpenStudy (amistre64):

3x^2 + 4x - 2 = mx - 5 3x^2 + 4x - 2 = mx - 5 -mx+5 -mx+5 --------------------- 3x^2+(4-m)x +3 = 0 looks to be the route sam was going

OpenStudy (amistre64):

-(4-m) +- sqrt((4-m)^2 - 4(3)(3)) ----------------------------- 2(3) m-4 +- sqrt(16+m^2-8m - 36) ----------------------------- 6 m-4 +- sqrt(m^2-8m - 20) ------------------------- 6